# Quantum algorithm for the gradient of a logarithm-determinant

### v1

@@ Abstract
The logarithm-determinant is a common quantity in many areas of physics and computer science. Derivatives of the logarithm-determinant compute physically relevant quantities in statistical physics models, quantum field theories, as well as the inverses of matrices. A multi-variable version of the quantum gradient algorithm is developed here to evaluate the derivative of the logarithm-determinant. From this, the inverse of a sparse-rank input operator may be determined efficiently. Measuring an expectation value of the quantum state–instead of all $N^{2}$ elements of the input operator–can be accomplished in $O(k\sigma)$ time in the idealized case for $k$ relevant eigenvectors of the input matrix. A factor $\sigma=\frac{1}{\varepsilon^{3}}$ or $-\frac{1}{\varepsilon^{2}}\log_{2}\varepsilon$ depends on the version of the quantum Fourier transform used for a precision $\varepsilon$. Practical implementation of the required operator will likely need $\log_{2}N$ overhead, giving an overall complexity of $O(k\sigma\log_{2}N)$. The method applies widely and converges super-linearly in $k$ when the condition number is high. For non-sparse-rank inputs, the algorithm can be evaluated provided that an equal superposition of eigenstates is provided. The output is given in $O(1)$ queries of oracle, which is given explicitly here and only relies on time-evolution operators that can be implemented with arbitrarily small error. The algorithm is envisioned for fully error-corrected quantum computers but may be implementable on near-term machines. We discuss how this algorithm can be used for kernel-based quantum machine-learning.
@@ I Introduction
This motivates us to look for a quantum algorithm that can generically compute the logarithm-determinant in a stable way and with fewer resources than the classical approach. Such a solution would have wide-ranging implications on the ability to compute quantities in many areas. [^2] Some review with mathematics is provided in Appendix A.
@@ II Quantum gradient of the logarithm-determinant
A similar proof follows from the trace-logarithm form that is equivalent to the logarithm-determinant (see Appendix A).
@@ II.1 Determination of eigensolutions
At the start of this algorithm, $k$ eigenvalues are required from an input $X$. We provide in Appendix B how to do this from a block-Lanczos technique [15, 16] and summarized in Box 1. In principle, block-Lanczos is not required and other methods may be used in place of it if they achieve sufficient accuracy of the eigenvectors. For the current paper, this technique is applied to demonstrate that the wavefunction preparation is not exponentially large [32].
@@ A Hadamard transform is applied onto one register
Figure 2: Quantum gradient algorithm for the derivatives of eigenvalues on eigenstates. The results of a block Lanczos calculation are used to construct unitaries $\Gamma^{(p)}$ for each of the $p$ eigenvalues to capture a sufficiently converged $S$ sparse-rank representation of the input matrix $X$. The operator $\Gamma^{(p)}$ will depend on a block Lanczos solution that was discussed in Sec. B. The remaining step is a quantum phase estimation with rotational gates controlled on the variations $\epsilon$. This produces a phase related to the eigenvalue, the desired result.
@@ II.3 Oracle: quantum gradient phase estimation
The oracle query for the quantum gradient algorithm can be phrased as a quantum phase estimation controlled on the second register in the QGA. This gives a scaling of QGPE like QPE of $O(m)\equiv O(-\log_{2}\varepsilon)$ for $m$ qubits giving precision of $\varepsilon$.
@@
we can determine that the expansion of the perturbing term $L\epsilon\Delta(i,j)/N$ follows as a series expansion of the term as seen in Eq. (17) (see Appendix C for the full form of a multi-variable series). Higher order terms in Eq. (19) therefore appear as $O(L^{2})$ corrections. The prefactor from Eq. (16) of $N/(ML)$ will cancel out the extraneous factor giving an overall term of $O(L)$. As $L$ is decreased, the error will vanish. In total, higher order terms will go as $O(L^{o-1})$ for order $o$ and so vanish faster than previous orders.
@@ II.4 Expectation values of the inverse
8:     QGPE onto wavefunction $|p\rangle$ controlled on $|\epsilon\rangle^{\otimes m}$ to produce the phase $e^{2\pi\epsilon i\delta E(x_{ij})}$triangleright Eq. (16)
9:     $|\delta E(x_{ij})\rangle\overset{\mathrm{QFT}}{\leftarrow}e^{2\pi\epsilon i\delta E(x_{ij})}|\epsilon\rangle^{\otimes m}$ and measure
@@ II.5 Complexity scaling
Classical algorithms that have pursued similar strategies [42, 43] have been reported to have $O(k^{2}N)$ scaling [44], meaning that the algorithm here is less by an exponential factor (a ratio of $\frac{\log_{2}N}{kN}$ for some precision) in the practical case.
As a comment on gate-depth for this algorithm, it is entirely dependent on the (controlled) gates that are applied in acting the operators of QGPE onto the quantum machine and their implementation [45].
@@ II.6 Discussion
The existence of this algorithm seems to match the expected optimal scaling by related works on the quantum singular value transformation [46, 47, 48]. The quantum algorithm is less than classical algorithms that we are aware of by an exponential factor [44] in the case of the practical scaling ($N\rightarrow\log_{2}N$).
This algorithm can be viewed as a dual algorithm to Ref. 49 in the sense that Ref. 49 focuses on the input wavefunction rather than the operator [50, 40, 51, 52, 53, 54, 32, 55]. The algorithm here is focused on treating the operator itself.
@@ III Applications to sparse-rank input: kernel-based machine learning
We note that at least $k$ relevant eigenvectors must be present for the algorithm to obtain an answer in general. We find the most useful complexity proof demonstrating this to be native to machine learning. It has been proven in Ref. 62 (with concurring results in at least Ref. 63 and Ref. 64) that the number of oracle queries for a quantum machine learning algorithm are equal to the number of training points on the classical computer (i.e., quantum computers store no more correlations than are provided to it; same for the classical computer). So, at least this much information must be provided in the wavefunction, else the proof from Ref. 62 would need modification and imply that quantum machine learning is not restricted by the probably approximately correct (PAC) model of learning at least [65, 66].
@@ IV Extensions for other state preparation assumptions
We note that this is related to the same wavefunction preparation assumptions as other works, namely Ref. 49 [53, 54, 32, 55]. If a suitable set of preparation assumptions can be imposed (potentially with a quantum memory), then we will explore that possibility in this section.
@@ IV.1 Preparation of an equal superposition of eigenstates
We do note the recent appearance of some methods to find excitation energies on the quantum computer that may be useful [68, 69, 16, 70].
@@ IV.5 Application of $\Sigma$-QGLD
This method is perhaps best applied when the eigenstates are known, such as for the uniform gas. In this case, we take a collection of fermions in the unit cell of an infinite system [72]. The external potential is set to cancel a divergence in the size-extensive exchange term (although, this can be taken to be zero for all practical purposes), and the eigenstates are considered to be plane waves $\exp(i\mathbf{k}\cdot\mathbf{r})$. In this case, we can prepare the set of eigenstates on the quantum computer through a unitary.
@@ V Conclusion
The goal was to show that the logarithm-determinant can be solved on the quantum computer. There is no claim about the ultimate performance with respect to all possible (un)known classical algorithms [53, 54, 32, 55].
@@ VI Acknowledgements
We thank Olivia Di Matteo, Sarah Huber, Michael O. Flynn, Arif Babul, and Emmanuel Fromager for useful comments. T.E.B. is forever indebted to David Poulin for the original inspiration and time that led to this work.
@@ VII Author Contributions
Appendix A Logarithm-determinant in physics
@@
Appendix B Random quantum block Lanczos recursion
@@
Lanczos methods have recently caught the imagination of the quantum computing development community. After the original proposal with block encodings in Ref. 15, there was a nearly immediate demonstration on a real quantum computer of a continued fraction in Ref. 80. The extension to block Lanczos techniques followed shortly after that [16], noting that the method should be more stable than the scalar Lanczos case based on error estimates. An in-depth summary of the method covered existing strategies in Ref. 81–reviewing in detail that the implementation on the quantum computer is efficient–with some discussion of error in Ref. 82. The problem has recently been applied with near-term quantum technologies to large spin chains [83] and lattice descriptions of quantum field theories [84]. Some other implementations of Lanczos methods use Trotter decompositions of a time evolution operator [85, 86, 87, 88]. Some other methods have investigated Davidson methods for excited states [89].
@@
In all, this represents the initial $b$ states for a block-size $b$ of the Lanczos algorithm. The initial states should be made to be orthogonal to each other. One easy way to create the states on the classical computer is to initialize an $N\times b$ sized matrix $\Psi_{0}$ randomly and then take a singular value decomposition of the matrix $\mathrm{svd}(\Psi_{0})=UDV^{\dagger}$ [90, *baker2019m] and then reassign $\Psi_{0}=UV^{\dagger}$. This will ensure orthogonality. In principle, the initial wavefunction can then be prepared on the quantum computer.
@@
Appendix C Multi-variable series expansion
@@
Appendix D Classical checks on small operators
@@
Derivatives in this section are taken with respect to finite difference approximations of the perturbed matrix (non-quantum implementation). The full quantum implementation, which matches the exact exact result roughly to $O(L)$, is detailed in Appendix E.
@@ D.1 $X=\sigma^{x}$
The quantum algorithm was implemented in a simulator in the DMRjulia library [90, *baker2019m, 92, 93, 94], and the result when using a single qubit register ($m=1$) matches the slope in Fig. 3.
D.2 $X=\sigma^{z}$
@@ D.3 $X=H$
Appendix E Implementation of the Quantum Gradient Phase Estimation
@@
E.1 Oracle query for the gradient
@@
E.2 Implementation of the quantum Fourier transform
@@
E.3 Example 1: Pauli X Gate
@@ E.4 Results
E.5 Example 2: Hadamard Gate
@@
E.6 Simulation details
We built a quantum simulator on top of a matrix product state in the DMRjulia library [90, *baker2019m, 92, 93, 94]. All basic algorithms were tested to ensure accuracy. Implementation details of this library will be detailed later, but the simulator is compiler-assumption-free.
Appendix F QGLD sampling

### v2

@@ Abstract
The logarithm-determinant is a common quantity in many areas of physics and computer science. Derivatives of the logarithm-determinant compute physically relevant quantities in statistical physics models, quantum field theories, as well as the inverses of matrices. A multi-variable version of the quantum gradient algorithm is developed here to evaluate the derivative of the logarithm-determinant. From this, the inverse of a sparse-rank input operator may be determined efficiently. Measuring an expectation value of the quantum state–instead of all $N^{2}$ elements of the input operator–can be accomplished in $O(k\sigma)$ time in the idealized case for $k$ relevant eigenvectors of the input matrix. A factor $\sigma=\frac{1}{\varepsilon^{3}}$ or $-\frac{1}{\varepsilon^{2}}\log_{2}\varepsilon$ depends on the version of the quantum Fourier transform used for a precision $\varepsilon$. Practical implementation of the required operator will likely need $\log_{2}N$ overhead, giving an overall complexity of $O(k\sigma\log_{2}N)$. The method applies widely and converges super-linearly in $k$ when the condition number is high. The best classical method we are aware of scales as $N$.
Given the same resource assumptions as other algorithms, such that an equal superposition of eigenvectors is available efficiently, the algorithm is evaluated in the practical case as $O(\sigma\log_{2}N)$. The output is given in $O(1)$ queries of oracle, which is given explicitly here and only relies on time-evolution operators that can be implemented with arbitrarily small error. The algorithm is envisioned for fully error-corrected quantum computers but may be implementable on near-term machines. We discuss how this algorithm can be used for kernel-based quantum machine-learning.
@@ I Introduction
This motivates us to look for a quantum algorithm that can generically compute the logarithm-determinant in a stable way and with fewer resources than the classical approach. Such a solution would have wide-ranging implications on the ability to compute quantities in many areas. [^2] Some review with mathematics is provided in Appendix B.
@@ II Quantum gradient of the logarithm-determinant
A similar proof follows from the trace-logarithm form that is equivalent to the logarithm-determinant (see Appendix B).
@@ II.1 Determination of eigensolutions
At the start of this algorithm, $k$ eigenvalues are required from an input $X$. We provide in Appendix C how to do this from a block-Lanczos technique [15, 16] and summarized in Box 1. In principle, block-Lanczos is not required and other methods may be used in place of it if they achieve sufficient accuracy of the eigenvectors. For the current paper, this technique is applied to demonstrate that the wavefunction preparation is not exponentially large [32].
@@ A Hadamard transform is applied onto one register
Figure 2: Quantum gradient algorithm for the derivatives of eigenvalues on eigenstates. The results of a block Lanczos calculation are used to construct unitaries $\Gamma^{(p)}$ for each of the $p$ eigenvalues to capture a sufficiently converged $S$ sparse-rank representation of the input matrix $X$. The operator $\Gamma^{(p)}$ will depend on a block Lanczos solution that was discussed in Sec. C. The remaining step is a quantum phase estimation with rotational gates controlled on the variations $\epsilon$. This produces a phase related to the eigenvalue, the desired result.
@@ II.3 Oracle: quantum gradient phase estimation
The oracle query for the quantum gradient algorithm can be phrased as a quantum phase estimation controlled on an equal superposition of bit strings. This gives a scaling of QGPE like QPE of $O(m)\equiv O(-\log_{2}\varepsilon)$ for $m$ qubits giving precision of $\varepsilon$.
@@
we can determine that the expansion of the perturbing term $L\epsilon\Delta(i,j)/N$ follows as a series expansion of the term as seen in Eq. (17) (see Appendix D for the full form of a multi-variable series). Higher order terms in Eq. (19) therefore appear as $O(L^{2})$ corrections. The prefactor from Eq. (16) of $N/(ML)$ will cancel out the extraneous factor giving an overall term of $O(L)$. As $L$ is decreased, the error will vanish. In total, higher order terms will go as $O(L^{o-1})$ for order $o$ and so vanish faster than previous orders.
@@ II.4 Expectation values of the inverse
8:     QGPE onto wavefunction $|p\rangle$ controlled on $|\epsilon\rangle^{\otimes m}$ to produce the phase $e^{2\pi\epsilon i\delta E_{p}(x_{ij})}$triangleright Eq. (16)
9:     $|\delta E_{p}(x_{ij})\rangle\overset{\mathrm{QFT}}{\leftarrow}e^{2\pi\epsilon i\delta E_{p}(x_{ij})}|\epsilon\rangle^{\otimes m}$ and measure or replacing $\delta E(x_{ij})\rightarrow\sum_{ij}\delta E_{p}(x_{ij})$ for all perturbed entries
@@ II.6 Discussion
The existence of this algorithm seems to match the expected optimal scaling by related works on the quantum singular value transformation [42, 43, 44].
The quantum algorithm is less than classical algorithms that we are aware of by an exponential factor [45] in the case of the practical scaling ($N\rightarrow\log_{2}N$). Classical algorithms that have pursued similar strategies [46, 47] have been reported to have $O(k^{2}N)$ scaling [45], meaning that the algorithm here is less by an exponential factor (a ratio of $\frac{\log_{2}N}{kN}$ for some precision) in the practical case.
This algorithm can be viewed as a dual algorithm to Ref. 48 in the sense that Ref. 48 focuses on the input wavefunction rather than the operator [49, 40, 50, 51, 52, 53, 32, 54, 55]. The algorithm here is focused on treating the operator itself. We can only speculate on exactly what operators will be best for this algorithm and so cannot comment on exactly how efficient the implementation will be in light of future progress. We do note that writing the operator in a format amenable to the quantum computer may require a different form for the QGPE; we merely followed Ref. 30’s form.
It has been noted before that similar algorithms [48] are highly dependent on their state preparation assumptions [44], and this dual algorithm is likely highly dependent on the ability to implement input time-evolution gate operators in an efficient way. However, for many cases of interest, this problem is a common feature of many algorithms, so improved strategies with implementing time-evolution gates is likely to occur in other areas and may enable more use applications here. We suggest that this feature, being common to many other algorithms, may have a clearer path to improvement over the wavefunction preparation techniques required for other algorithms [44, 54]. The gate depth will be entirely dependent on the (controlled) gates that are applied in acting the operators of QGPE onto the quantum machine and their implementation [56].
We further note that Ref. 48 makes use of a sparsity assumption for the input operator. We can in principle do so here, but may have to compute many derivatives regardless of the sparsity assumption. A block form of the input operator would be a natural equivalent here.
@@ III Applications to sparse-rank input: kernel-based machine learning
We note that at least $k$ relevant eigenvectors must be present for the algorithm to obtain an answer in general. We find the most useful complexity proof demonstrating this to be native to machine learning. It has been proven in Ref. 63 (with concurring results in at least Ref. 64 and Ref. 65) that the number of oracle queries for a quantum machine learning algorithm are equal to the number of training points on the classical computer (i.e., quantum computers store no more correlations than are provided to it; same for the classical computer). So, at least this much information must be provided in the wavefunction, else the proof from Ref. 63 would need modification and imply that quantum machine learning is not restricted by the probably approximately correct (PAC) model of learning at least [66, 67].
@@ IV Extensions for other state preparation assumptions
We note that this is related to the same wavefunction preparation assumptions as other works, namely Ref. 48 [52, 53, 32, 54]. If a suitable set of preparation assumptions can be imposed (potentially with a quantum memory), then we will explore that possibility in this section.
@@ IV.5 Application of $\Sigma$-QGLD
This method is perhaps best applied when the eigenstates are known, such as for the uniform gas. In this case, we take a collection of fermions in the unit cell of an infinite system [70]. The external potential is set to cancel a divergence in the size-extensive exchange term (although, this can be taken to be zero for all practical purposes), and the eigenstates are considered to be plane waves $\exp(i\mathbf{k}\cdot\mathbf{r})$. This has been explored on the quantum computer before [71, 72, 73, 74]. In this case, we can prepare the set of eigenstates on the quantum computer through a unitary.
@@ V Conclusion
The goal was to show that the logarithm-determinant can be solved on the quantum computer. There is no claim about the ultimate performance with respect to all possible unknown classical algorithms [52, 53, 32, 54].
@@ VI Acknowledgements
We thank Olivia Di Matteo, Nishanth Baskaran, Sarah Huber, Michael O. Flynn, Arif Babul, and Emmanuel Fromager for useful comments. T.E.B. is forever indebted to David Poulin for the original inspiration and time that led to this work.
@@
Some figures were prepared with the Quantikz package [77].
@@ VII Author Contributions
Appendix A Negative determinants and eigenvalues
There is a question of the applicability of Eq. (1) when the determinant of a matrix is negative. The natural logarithm is defined only in the complex plane by $\ln z=\ln|z|+i\theta$ where $\theta$ is chosen as a branch cut arbitrarily. Often, this is chosen to be along the negative real axis, but this can be chosen elsewhere. In our application here, we can avoid the appearance of the imaginary number by noting that the derivative in Eq. (1) will cause the imaginary number to be zero on evaluation.
Even if the negative determinant appears, the evaluation from Eq. (2) still holds. As an example, considering the case of $X=\sigma^{z}$, we have
$\sigma^{z}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)$ (38)
with eigenvectors $|0\rangle=(1,0)^{T}$ and $|1\rangle=(0,1)^{T}$. The evaluation of $\langle p|e^{2\pi i\frac{N}{ML}(Z+\frac{L\epsilon}{N}\Delta)}|p\rangle$. will still produce the correct result despite having negative eigenvalues. The overall sum is zero since $(\sigma^{z})^{-1}=\sigma^{z}$ leading to a full sum being 0 over all $p$. However, we only find positive derivatives of the derivatives of the eigenvalues. We must have $E_{p}=\pm 1$ in the alternative form of Eq. (2). Thus, a negative determinant can be handled by the predominant form that we use from Eq. (2).
Appendix B Logarithm-determinant in physics
@@
Appendix C Random quantum block Lanczos recursion
@@
Lanczos methods have recently caught the imagination of the quantum computing development community. After the original proposal with block encodings in Ref. 15, there was a nearly immediate demonstration on a real quantum computer of a continued fraction in Ref. 83. The extension to block Lanczos techniques followed shortly after that [16], noting that the method should be more stable than the scalar Lanczos case based on error estimates. An in-depth summary of the method covered existing strategies in Ref. 84–reviewing in detail that the implementation on the quantum computer is efficient–with some discussion of error in Ref. 85. The problem has recently been applied with near-term quantum technologies to large spin chains [86] and lattice descriptions of quantum field theories [87]. Some other implementations of Lanczos methods use Trotter decompositions of a time evolution operator [88, 89, 90, 91]. Some other methods have investigated Davidson methods for excited states [92].
@@
In all, this represents the initial $b$ states for a block-size $b$ of the Lanczos algorithm. The initial states should be made to be orthogonal to each other. One easy way to create the states on the classical computer is to initialize an $N\times b$ sized matrix $\Psi_{0}$ randomly and then take a singular value decomposition of the matrix $\mathrm{svd}(\Psi_{0})=UDV^{\dagger}$ [93, *baker2019m] and then reassign $\Psi_{0}=UV^{\dagger}$. This will ensure orthogonality. In principle, the initial wavefunction can then be prepared on the quantum computer.
@@
Appendix D Multi-variable series expansion
@@
Appendix E Classical checks on small operators
@@
Derivatives in this section are taken with respect to finite difference approximations of the perturbed matrix (non-quantum implementation). The full quantum implementation, which matches the exact exact result roughly to $O(L)$, is detailed in Appendix F.
@@ E.1 $X=\sigma^{x}$
The quantum algorithm was implemented in a simulator in the DMRjulia library [93, *baker2019m, 95, 96, 97], and the result when using a single qubit register ($m=1$) matches the slope in Fig. 3.
E.2 $X=\sigma^{z}$
@@ E.3 $X=H$
Appendix F Implementation of the Quantum Gradient Phase Estimation
@@
F.1 Oracle query for the gradient
@@
F.2 Implementation of the quantum Fourier transform
@@
F.3 Example 1: Pauli X Gate
@@ F.4 Results
F.5 Example 2: Hadamard Gate
@@
F.6 Simulation details
We built a quantum simulator on top of a matrix product state in the DMRjulia library [93, *baker2019m, 95, 96, 97]. All basic algorithms were tested to ensure accuracy. Implementation details of this library will be detailed later, but the simulator is compiler-assumption-free.
Appendix G QGLD sampling